Given the pair of linear equations in two variables.
Find x and y intercepts for each line. Apply elimination or substitution to find the coordinates of the point of interception.
Problem:
9x - 4y = 8
5x + 3y = 41

Answer by JBarnum(2146)

Elimination version:
9x - 4y = 8 multiply this line by 3
5x + 3y = 41 multiply this line by 4
21x - 12y = 24 add the lines together
20x + 12y = 164
41x =188
4.5854=x
5(4.5854) +3y =41
17.1429 +3y =41
3y= 23.8571
7.9524=y